# Magnetic, Kinetic, and Transition regime: Spatially-segregated structure of compressive MHD turbulence

### v1

@@ Abstract
Turbulence is a complex physical process that emerges in multiple areas of modern physics, and in ionized environments such as interstellar gas, the magnetic field can be dynamically important. However, the exact function of the magnetic field in the ionized gas remains unclear. We use the ${\cal M}_{\rm A}=\sqrt{E_{\rm k}/E_{B}}$ to describe the importance of the magnetic field measured to the turbulent motion, and reveal diverse ways of mutual interaction. At low ${\cal M}_{\rm A}$ (magnetic regime), the magnetic field is well-described as force-free. Despite the strong magnetic field, the motion of gas does not stay aligned with the magnetic field. At the regime of intermediate ${\cal M}_{\rm A}$ (magnetic-kinetic transition regime), the velocity field and the magnetic field exhibit the highest degree of alignment, which is likely the result of a rapid relaxation. At high ${\cal M}_{\rm A}$ (kinetic regime), both the magnetic field and the velocity field are irregular, with no alignment. We find observational counterparts to these regimes in observations of interstellar gas. The results highlight the diverse behavior of gas in MHD turbulence and guide future interpretations of the role of the magnetic field in astrophysical observations.
@@ 1 Introduction
Turbulence is a complex process that has puzzled scientists since the age of Leonardo da Vinci, and the capability of turbulence in controlling the evolution of interstellar gas has been known for decades (Mac Low & Klessen, 2004). Understanding the role of the magnetic field in compressible turbulence can be crucial for our understanding of turbulence and for interpreting astrophysical observations. Turbulence is a complex, multi-scale process Frisch (1995) best described using scaling relations Kolmogorov (1941). Past studies of compressible magnetohydrodynamics (MHD) turbulence have followed this tradition where describing the statistic properties of the region has become a priority Burkhart et al. (2020). Others have astrophysical applications in mind and have focused on the global quantities extracted from the simulation box (Padoan & Nordlund, 2011).
The alignment between vector quantities such as the magnetic field $\vec{B}$, the velocity $\vec{v}$, and the current $J$ offers insight into the behavior of the magnetized fluids. In astrophysical research, the alignment between $\vec{B}$ and $\vec{v}$ is assumed to be an indicator of the magnetic field’s ability to affect the motion of the gas. The alignment between $\vec{B}$ and $\vec{v}$ in the strongly magnetized regime is a fact often taken for granted, and this picture is the foundation for understanding phenomena such as wind from disk-star systems Pudritz et al. (2007), where the picture of beads on a wire have been widely accepted. In this picture, field lines of a magnetic field behave like rigid wires, which guide the motion of the gas. However, it is unclear to what extent can we trust this picture. On the other hand, Matthaeus et al. (2008) have shown that the rapid alignment between $\vec{v}$ and $\vec{B}$ is the result of a rapid relaxation process, where the magnetic field does not need to dominant.
@@ 2 Data and method
We use numerical simulations of MHD equations performed using the Enzo code (Collins et al., 2010; Burkhart et al., 2015) with the constrained transport turned on. The simulation conducted in this study analyzed the impact of self-gravity and magnetic fields on supersonic turbulence in isothermal molecular clouds, using high-resolution simulations and adaptive mesh refinement techniques (Collins et al., 2012; Burkhart et al., 2015, 2020). The simulation we use has initial $\beta\approx E_{\rm k}/E_{\rm B}=0.2$. However, as the simulation proceeds, different subregions have different $\beta$ and ${\cal M}_{\rm A}$.
@@
Figure 2: Alignment between magnetic field and kinetic motion. The panel shows the distribution of magnetic energy density $E_{B}$ (blue) and kinetic energy density $E_{\rm k}$ (red). Orange and blue vectors present the orientation of the magnetic field and velocity. The contours show the -0.5 (cyan) and 0. 5 (orange) of log10 ${\cal M}_{\rm A}$. Locations of gas belonging to different regimes are outlined in the plot at the upper left corner.
Figure 3: Alignment between magnetic field and current. The panel shows the distribution of magnetic field strength $|\vec{B}|$ (blue) and current density $|\vec{J}|$ (red). Orange and blue vectors present the orientation of the magnetic field and velocity. The contours show the log10 ${\cal M}_{\rm A}$= -0.5 (cyan) and 0. 5 (orange) respectively. Locations of gas belonging to different regimes are outlined in the plot at the upper left corner.
@@ 3 Three regimes of MHD Turbulence
Based on the Alfven Mach number ${\cal M}_{\rm A}$, we divide the simulation box into three regimes: magnetic regime (${\cal M}_{\rm A}<0.3$), B-k transition ($30.3<{\cal M}_{\rm A}<$ 3), and kinetic regime (${\cal M}_{\rm A}>3$), and study the alignment between the magnetic field $\vec{B}$, velocity $\vec{v}$, and current $\vec{J}$ at different ${\cal M}_{\rm A}$. The relationship between the magnetic field-velocity angle $\theta_{\vec{B}-\vec{v}}$, the magnetic field-current angle $\theta_{\vec{B}-\vec{J}}$ and the Alfven Mach number is shown in Fig. 1. When plotting these distributions, we subtract the distribution expected if the vectors are randomly oriented, and focus on the additional alignment caused by physics (see Sec. A).
@@
  • The magnetic regime (${\cal M}_{\rm A}<0.3$): the magnetic energy is far above the local kinetic energy, where the $\vec{B}$ and $\vec{v}$ do not stay aligned, yet the $\vec{B}$ and $\vec{J}$ aligned.
  • The transition regime (B-k transition, $30.3<{\cal M}_{\rm A}<$ 3): the magnetic energy and kinetic energy have similar densities. the $\vec{B}$ and $\vec{v}$ stay aligned, and the $\vec{B}$ and $\vec{J}$ evolve from aligned to not aligned as ${\cal M}_{\rm A}$ increases.
@@
which are plotted in Figs. 2, 3 and 4, and are summarized in Fig. 4. Some addition slices to the simulation box can be found in Sec. B
@@ 3.2 B-k Transition
At $0.3<\cal{M}_{\rm A}$ $<3$, where the magnetic energy and the kinetic energy are similar, we find a transition regime characterized by a breakdown of the alignment between $\vec{B}$ and $J$, and a strong alignment between $\vec{B}$ and $\vec{v}$. The breakdown of the $\vec{B}$-$J$ alignment results from a decrease in the importance of the magnetic field. However, the alignment between $\vec{B}$ and $\vec{v}$ deserves further discussions.
We find the $\vec{B}$ and $\vec{v}$ only occur in the B-k transition regime, where the magnetic and kinetic energy have similar densities. This finding challenges the common understanding that the alignment between $\vec{B}$ and $\vec{v}$ indicates the dominance of the magnetic field. This alignment has been found by Matthaeus et al. (2008), which results from a “rapid and robust relaxation process in turbulent flows”. Our findings support their conclusion. However, different from the claim that “ the alignment of the velocity and magnetic field fluctuations occurs rapidly in magnetohydrodynamics for a variety of parameters”, in our case, this alignment only occurs at the transition regime with moderate $\cal{M}_{\rm A}$. In the B-k transition regime, the $vec{B}$ and $\vec{J}$ are still moderately aligned with $\theta\approx 15^{\circ}$.
@@ 3.3 Kinetic Regime
At the kinetic regime with high kinetic energy (high ${\cal M}_{\rm A}$), the $\vec{B}-\vec{v}$ not aligned, and the $\vec{B}-\vec{J}$ not aligned. The lack of this alignment is the result of the weak magnetic field.
In astronomical observations, this corresponds to dense, collapsing regions with strong turbulence, such as the Cyg N44 in Dr21 (Ching et al., 2018).
@@ 4 Conclusions
We investigate the effect of a magnetic field on supersonic turbulence. By evaluating the Alfven Mach number $\cal{M}_{\rm A}$ at different locations in a turbulence box and studying the alignment between the magnetic field $\vec{B}$, the velocity $\vec{v}$ and the current $\vec{J}$, we reveal the different behavior of the system under various conditions. These regimes include
@@
  • The transition regime (B-k transition, $30.3<{\cal M}_{\rm A}<$ 3): the magnetic energy and kinetic energy have similar densities. $\vec{B}$ and $\vec{v}$ are. aligned, and the $\vec{B}$ and $\vec{J}$ evolve from aligned to not aligned as ${\cal M}_{\rm A}$ increases. We also observe alignments between the magnetic field and the gas motion. However, this alignment between $\vec{B}$ and $\vec{v}$ does not imp a strong magnetic field but a rapid relaxation process in turbulent flows.
@@
They are summarized in 4. Since there is a correlation between the Alfven Mach number and the gas density (Fig. 6), the magnetic regime exists in the lower-density part and the kinetic regime in the higher-density part. The transition regime is an intermediate state between the two. Using observational data, we find cases that support the existence of these regimes.
@@
We reveal various regimes where the fluid behaves differently under different conditions. To our knowledge, this is the first time these different regimes are clearly outlined. The alignment between these quantities has been studied in a recent paper Beattie et al. (2024) where the authors reported strong alignments in the strongly magnetized regime and some scale-dependence alignment behavior. We are revealing a much clearer picture with the Alfven Mach number as the only parameter dictating the behavior of the fluid and the alignments.
Figure 5: Summary of three regimes of compressive MHD turbulence. The Magnetic regime corresponds to observations in Heyer et al. (2016); Tritsis & Tassis (2016); Beattie et al. (2020); Skalidis et al. (2022). The B-k transition regime corresponds to observations in Hu et al. (2021); Zhao et al. (2022); Tram et al. (2023). The kinetic regime corresponds to observations toward Cyg N44 in Dr21 (Ching et al., 2018).
Figure 6: Relation between the Alfven Mach number $\cal{M}_{\rm A}$ and gas density $\rho$.
Appendix A Removing the projection effect in angle distributions
The probability density function, p($\theta$), of the angle between two vectors of random distribution is distributed in an N-dimensional space as:
$p(\theta)=\frac{\Gamma(\frac{n}{2})}{\Gamma(\frac{n-1}{2})}\frac{{\rm sin}^{n-2}(\theta)}{\sqrt{\pi}}$ (A1)
where n is numbers of dimensions, and $\theta$ is the angle between two vectors of random distribution. In our alignment analysis, we study the angle between qualities measured in 3D. When $n=3$, we have
$p(\theta)=\frac{1}{2}{\rm sin}\theta\;,$ (A2)
which two randomly-selected vectors have a alignment angle clustered at around $\theta=45^{\circ}$, caused by the projection effect. To remove the projection effect, we use the probability density function p($\theta$) to weight the angle distribution:
$I(\theta_{\rm corrected})=I(\theta_{\rm original})/p(\theta)$ (A3)
where $I(\theta_{\rm original})$ represents the original angle distribution and $I(\theta_{\rm corrected})$ represents the corrected angle distribution, with the projected effects removed.
Appendix B 2D slice of magnetic field, velocity and current
Fig. 7 shows the distribution of density $\rho$, velocity field $v$, magnetic field $B$, and current field $J$ in X-Y plane, which is a slice of 3D space.

### v2

@@ Abstract
Turbulence is a complex physical process prevalent in modern physics, particularly in ionized environments like interstellar gas, where magnetic fields play a dynamic role. However, the precise influence of magnetic fields in such settings remains unclear. We employ the Alfvén Mach number, ${\cal M}_{\rm A}=\sqrt{E_{\rm k}/E_{B}}$, to gauge the magnetic field’s significance relative to turbulent motion, uncovering diverse interaction patterns. In the low-${\cal M}_{\rm A}$ magnetic regime, the field is force-free, yet gas motion does not align with it. At intermediate ${\cal M}_{\rm A}$ (magnetic-kinetic transition regime), velocity and magnetic fields show peak alignment, likely due to rapid relaxation. In the high-${\cal M}_{\rm A}$ kinetic regime, both fields are irregular and unaligned. These regimes find observational counterparts in interstellar gas, highlighting the multifaceted nature of MHD turbulence and aiding future astrophysical interpretations.
keywords:
Magnetic fields – Magnetohydrodynamics (MHD) – ISM: clouds – ISM: kinematics and dynamics – ISM: magnetic fields
@@ 1 Introduction
Turbulence is a complex process that has puzzled scientists since the age of Leonardo da Vinci, and the capability of turbulence in controlling the evolution of interstellar gas has been known for decades (Mac Low & Klessen, 2004). Understanding the role of the magnetic field in compressible turbulence can be crucial for our understanding of turbulence and for interpreting astrophysical observations. Turbulence is a complex, multi-scale process (Frisch, 1995) best described using scaling relations (Kolmogorov, 1941). Past studies of compressible magnetohydrodynamics (MHD) turbulence have followed this tradition where describing the statistic properties of the region has become a priority (Burkhart et al., 2020). Others have astrophysical applications in mind and have focused on the global quantities extracted from the simulation box (Padoan & Nordlund, 2011).
The alignment between vector quantities such as the magnetic field $\vec{B}$, the velocity $\vec{v}$, and the current $J$ offers insight into the behavior of the magnetized fluids. In astrophysical research, the alignment between $\vec{B}$ and $\vec{v}$ is assumed to be an indicator of the magnetic field’s ability to affect the motion of the gas. The alignment between $\vec{B}$ and $\vec{v}$ in the strongly magnetized regime is a fact often taken for granted, and this picture is the foundation for understanding phenomena such as wind from disk-star systems (Pudritz et al., 2007), where the picture of beads on a wire have been widely accepted. In this picture, field lines of a magnetic field behave like rigid wires, which guide the motion of the gas. However, it is unclear to what extent can we trust this picture. On the other hand, alignment between $\vec{B}$ and $\vec{v}$ has been proposed by (Boldyrev, 2006). However, the alignment is analyzed in the Fourier space, where the spatial structure and possible segregation effect can not be analyzed. Matthaeus et al. (2008) have shown that the rapid alignment between $\vec{v}$ and $\vec{B}$ is the result of a rapid relaxation process, caused by an interplay between pressure gradient and the magnetic field when their strengths are comparable.
@@ 2 Data and method
We use numerical simulations of MHD equations performed using the Enzo code (Collins et al., 2010; Burkhart et al., 2015) with the constrained transport turned on. The simulation conducted in this study analyzed the impact of self-gravity and magnetic fields on supersonic turbulence in isothermal molecular clouds, using high-resolution simulations and adaptive mesh refinement techniques (Collins et al., 2012; Burkhart et al., 2015).They are available online https://www.mhdturbulence.com/, as presented in Burkhart et al. (2020). This simulation provides a 2563 cubes at super-sonic state ($M_{S,0}$ = 9) at the resolution of 0.018 pc. The simulation we use has initial $\beta_{0}\approx 0.2$. However, as the simulation proceeds, different subregions have different $\beta_{0}$ and ${\cal M}_{\rm A}$. We choose a snapshot at $t=0.9t_{\rm ff}$ upon which our analyses are performed. We also investigated other snapshots and initial conditions, and can confirm that the results, which links the behavior of the fluid with the Alfvenic Mach number measured at the microscopic scale are robust. The major different between those simulations reflects on the volume filling factor of the different regimes.
@@
Figure 2: Alignment between magnetic field and kinetic motion. The panel shows the distribution of magnetic energy density $E_{B}$ (blue) and kinetic energy density $E_{\rm k}$ (red). Orange and blue vectors present the orientation of the magnetic field and velocity. The contours show the -0.5 (darkred) and 0. 5 (orange) of log10 ${\cal M}_{\rm A}$. Locations of gas belonging to different regimes are outlined in the plot at the upper left corner.
Figure 3: Alignment between magnetic field and current. The panel shows the distribution of magnetic field strength $|\vec{B}|$ (blue) and current density $|\vec{J}|$ (green). Orange and blue vectors present the orientation of the magnetic field and velocity. The contours show the log10 ${\cal M}_{\rm A}$= -0.5 (darkred) and 0.5 (orange) respectively. Locations of gas belonging to different regimes are outlined in the plot at the upper left corner.
@@
Figure 5: Distribution of the offset angle $\theta_{\vec{B}-\vec{J}}$ and $\theta_{\vec{B}-\vec{v}}$ in three regimes.
Figure 7: A graphical summary of different regimes.
@@ 3 Three regimes of MHD Turbulence
Based on the Alfven Mach number ${\cal M}_{\rm A}$, we divide the simulation box into three regimes: magnetic regime (${\cal M}_{\rm A}<0.3$), B-k transition ($0.3<{\cal M}_{\rm A}<3$), and kinetic regime (${\cal M}_{\rm A}>3$), and study the alignment between the magnetic field $\vec{B}$, velocity $\vec{v}$, and current $\vec{J}$ at different ${\cal M}_{\rm A}$. The relationship between the magnetic field-velocity angle $\theta_{\vec{B}-\vec{v}}$, the magnetic field-current angle $\theta_{\vec{B}-\vec{J}}$ and the Alfven Mach number is shown in Fig. 1. When plotting these distributions, we divide our distribution by the distribution expected if the vectors are randomly oriented, and focus on the additional alignment caused by physics  [^1] The probability density function, p($\theta$), of the angle between two vectors of random distribution is distributed in an N-dimensional space as: $p(\theta)=\frac{\Gamma(\frac{n}{2})}{\Gamma(\frac{n-1}{2})}\frac{{\rm sin}^{n-2}(\theta)}{\sqrt{\pi}}$ (3) where n is numbers of dimensions, and $\theta$ is the angle between two vectors of random distribution. In our alignment analysis, we study the angle between qualities measured in 3D. When $n=3$, we have $p(\theta)=\frac{1}{2}{\rm sin}\theta\;,$ (4) which two randomly-selected vectors have a alignment angle clustered at around $\theta=45^{\circ}$, caused by the projection effect. To remove the projection effect, we use the probability density function p($\theta$) to weight the angle distribution: $I(\theta_{\rm corrected})=I(\theta_{\rm original})/p(\theta)$ (5) where $I(\theta_{\rm original})$ represents the original angle distribution and $I(\theta_{\rm corrected})$ represents the corrected angle distribution, with the projected effects removed. .
@@
  • The magnetic regime (${\cal M}_{\rm A}<0.3$): the magnetic energy is far above the local kinetic energy, where the $\vec{B}$ and $\vec{v}$ do not stay aligned, yet the $\vec{B}$ and $\vec{J}$ are aligned.
  • The transition regime (B-k transition, $0.3<{\cal M}_{\rm A}<3$): the magnetic energy and kinetic energy have similar densities. the $\vec{B}$ and $\vec{v}$ stay aligned, and the $\vec{B}$ and $\vec{J}$ evolve from aligned to not aligned as ${\cal M}_{\rm A}$ increases.
@@
which are plotted in Figs. 2, 3 and 4, and are summarized in Fig. LABEL:regimes. Some addition slices to the simulation box can be found in Fig. 6. We acknowledge that there is not clear boundary between these regimes. However, after dividing the simulation box into these three regimes, the alignment angle between the magnetic field $\vec{B}$, velocity $\vec{v}$, and current $\vec{J}$ do have distinct distributions (Fig. 5). 2D slice of magnetic field, velocity and current can be found in Fig. 6.
@@ 3.2 B-k Transition
At $0.3<\cal{M}_{\rm A}$ <3<3, where the magnetic energy and the kinetic energy are similar, we find a transition regime characterized by a breakdown of the alignment between $\vec{B}$ and $J$, and a strong alignment between $\vec{B}$ and $\vec{v}$. The breakdown of the $\vec{B}$-$J$ alignment results from a decrease in the importance of the magnetic field. However, the alignment between $\vec{B}$ and $\vec{v}$ deserves further discussions.
We find the $\vec{B}$ and $\vec{v}$ can achieve alignment in the B-k transition regime, where the magnetic and kinetic energy have similar densities. This finding challenges the common understanding that the alignment between $\vec{B}$ and $\vec{v}$ indicates the dominance of the magnetic field. This alignment has been found by Matthaeus et al. (2008), which results from a “rapid and robust relaxation process in turbulent flows”. Our findings support their conclusion. However, different from the claim that “ the alignment of the velocity and magnetic field fluctuations occurs rapidly in magnetohydrodynamics for a variety of parameters”, in our case, this alignment only occurs at the transition regime with moderate $\cal{M}_{\rm A}$. The fact that this alignment can occur only at a particular regime of of $\cal{M}_{\rm A}$ agrees with the findings of Kritsuk et al. (2017), where our critical $\cal{M}_{\rm A}$ should correspond to their critical density. In the B-k transition regime, the $\vec{B}$ and $\vec{J}$ are moderately aligned with $\theta\approx 15^{\circ}$.
@@ 3.3 Kinetic Regime
At the kinetic regime with high kinetic energy (high ${\cal M}_{\rm A}$), the $\vec{B}-\vec{v}$ not aligned, and the $\vec{B}-\vec{J}$ not aligned. The lack of this alignment is the result of the weak magnetic field. In astronomical observations, this corresponds to dense, collapsing regions with strong turbulence, such as the Cyg N44 in Dr21 (Ching et al., 2018).
@@ 4 Conclusions
We investigate the effect of a magnetic field on supersonic turbulence. By evaluating the Alfven Mach number $\cal{M}_{\rm A}$ at different locations in a turbulence box and studying the alignment between the magnetic field $\vec{B}$, the velocity $\vec{v}$ and the current $\vec{J}$, we reveal the different behavior of the system under various conditions. These regimes include:
@@
  • The transition regime (B-k transition, $0.3<{\cal M}_{\rm A}<3$): the magnetic energy and kinetic energy have similar densities. $\vec{B}$ and $\vec{v}$ are aligned, and the $\vec{B}$ and $\vec{J}$ are also aligned. We note this alignment between $\vec{B}$ and $\vec{v}$ does not necessarily imply a strong, dominant magnetic field but can also result from a rapid relaxation process in turbulent flows.
@@
They are summarized in 4. Since there is a correlation between the Alfven Mach number and the gas density e.g. (Zhao et al., 2024), the magnetic regime exists in the lower-density part and the kinetic regime in the higher-density part. The transition regime is an intermediate state between the two. Using observational data, we find cases that support the existence of these regimes.
@@
We reveal various regimes where the fluid behaves differently under different conditions. To our knowledge, this is the first time these different regimes are clearly outlined. The alignment between these quantities has been studied in a recent paper (Beattie et al., 2024) where the authors reported strong alignments in the strongly magnetized regime and some scale-dependence alignment behavior. We are revealing a much clearer picture with the Alfven Mach number as the only parameter dictating the behavior of the fluid and the alignments.
Data availability
The data used in this work is from the MHD simulation of Collins et al. (2012); Burkhart et al. (2015), which is publicly available at https://www.mhdturbulence.com/ presented in Burkhart et al. (2020).
Acknowledgements
GXL acknowledges support from NSFC grant No. 12273032 and 12033005. We thank Blakesley Burkhart for discussions on the maching-learning paper and contributors of https://www.mhdturbulence.com/ for making the data publicly available.